### v2

# Covariant dynamics from static spherically symmetric geometries

Abstract
We show that any one-parameter family of static, spherically symmetric spacetime metrics can be embedded as the unique vacuum solution of a generally covariant theory of pure gravity. These theories, involving at most second-order derivatives of metrics, are uniquely determined—up to two arbitrary functions that leave the metric unchanged—by a framework that resolves the problem of restoring general covariance in the Hamiltonian formulation. This result solves the inverse problem of Birkhoff’s theorem in full generality and establishes a systematic approach for reconstructing covariant dynamics from a given geometry. The framework applies broadly to quantum-gravity-inspired metrics, including regular black hole models, and provides a new tool for probing their dynamical and observational foundations.
I Introduction
Recent advances in observational techniques Abbott et al. (2016, 2017); Akiyama et al. (2019, 2022) have made it increasingly feasible to probe static, spherically symmetric spacetimes that deviate from the Schwarzschild solution, the unique spherically symmetric vacuum solution of general relativity (GR) guaranteed by Birkhoff’s theorem. Against this backdrop, the inverse problem of Birkhoff’s theorem arises naturally: how can one construct a self-consistent, covariant field-theoretic origin for novel static, spherically symmetric spacetimes beyond GR?
In addition to observational relevance, exploring metrics that deviate from GR is theoretically motivated by addressing fundamental issues like spacetime singularities. One common approach to generating such metrics involves ad hoc modifications of known black hole solutions Bardeen (1968); Bronnikov (2001); Hayward (2006); Ansoldi (2008); Cañate and Perez Bergliaffa (2023); Lan et al. (2023). The resulting geometries, known as regular black holes, are typically realized as solutions of GR coupled to exotic matter fields Ayon-Beato and Garcia (1998, 2000); Bronnikov and Fabris (2006); Bronnikov and Walia (2022); Chew et al. (2023); Li and Lü (2024). However, this strategy is widely viewed as unsatisfactory, not only because it requires exotic matter, but also because the underlying theories generally continue to admit singular solutions Huang and Rao (2025); Rao and Huang (2025). Moreover, the absence of a dynamical framework that naturally predicts regular black holes has hindered systematic studies of their formation and stability, a topic that has recently attracted growing interest Bueno et al. (2025a); Harada et al. (2025); Vertogradov et al. (2025); Bueno et al. (2025b).
Another approach beyond GR seeks to incorporate quantum corrections to classical black hole solutions, aiming to resolve singularities through fundamentally modified gravitational dynamics. One strategy within this framework involves gluing a static, spherically symmetric exterior metric to a singularity-free interior model describing dust collapse in a quantum gravity (QG) framework Bambi et al. (2013); Lewandowski et al. (2023); Bonanno et al. (2024). In such constructions, the exterior is typically determined by ad hoc junction conditions, while the absence of a consistent dynamical description for the full spacetime remains a key limitation. A more systematic strategy, that could overcome the limitations of these junction-based models, is to directly quantize the spherically symmetric sector of GR Modesto (2010); Gambini and Pullin (2008); Corichi and Singh (2016); Ashtekar et al. (2018); Husain et al. (2022); Gambini et al. (2023); Giesel et al. (2023); Ashtekar et al. (2023); Belfaqih et al. (2024). However, most of these models, as semiclassical theories derived from canonical approaches to QG, are formulated in the Hamiltonian framework, which necessarily requires a $3+1$ decomposition of spacetime. This decomposition, which splits spacetime into distinct space and time components, renders the preservation of general covariance highly nontrivial and thus poses a central challenge for such approaches Bojowald (2019); Gambini et al. (2022a); Bojowald (2022); Gambini et al. (2022b); Han and Liu (2024); Alonso-Bardaji and Brizuela (2024); Bojowald and Duque (2024); Alonso-Bardaji and Brizuela (2024); Zhang et al. (2025, 2024); Belfaqih et al. (2024); Yang et al. (2025). This challenge motivates previous works Zhang et al. (2025, 2024), conducted with the participation of a member of the present collaboration, which systematically address the problem of preserving general covariance within the Hamiltonian framework.
To set the stage, we recall the basic ingredients of the Hamiltonian framework for spherically symmetric GR on a 4-dimensional manifold $\mathcal{M}_{2}\times\mathbb{S}^{2}$. Here $\mathbb{S}^{2}$ denotes the 2-sphere, and $\mathcal{M}_{2}$ is an arbitrary 2-dimensional Lorentzian manifold. Let $(t,x)$ be a coordinate system defined in $\mathcal{M}_{2}$. In this coordinate system, the chart domain is foliated into constant-$t$ slices $\Sigma_{t}$, each diffeomorphic to some manifold $\Sigma$. Due to the spherical symmetry, the phase space consists of the canonical pairs $(K_{I},E^{I})$ with $I=1,2$, where, in the classical theory, $E^{I}$ encodes information about the spatial metric, while $K_{I}$ relates to the extrinsic curvature. The nonvanishing Poisson brackets of the canonical pairs read $\{K_{1}(x),E^{1}(y)\}=2\delta(x,y)$ and $\{K_{2}(x),E^{2}(y)\}=\delta(x,y),$ where the geometrized units with $G=1=c$ are applied. In classical GR, the 3+1 decomposition splits the Einstein equations into the Hamiltonian constraint, the diffeomorphism constraint, and the evolution equations. Since the evolution equations are equivalent to the Hamilton’s equations generated by a linear combination of the constraints, the essence of the dynamics lies in these two constraints.
In the present paper, we generalize all these classical structures to construct effective models that, for instance, could incorporate QG effects. In particular, following Zhang et al. (2025, 2024), we assume that the diffeomorphism constraint $H_{x}$ retains its classical form $H_{x}=E^{2}\partial_{x}K_{2}-K_{1}\partial_{x}E^{1}/2$ which generates diffeomorphism transformations on $\Sigma$. By contrast, the Hamiltonian constraint is expected to be modified into an effective one, denoted by $H_{\rm eff}$}, in order to account for QG effects. To preserve general covariance, the form of $H_{\rm eff}$} must be restricted, since certain choices are known to break covariance Kelly et al. (2020). In Zhang et al. (2025, 2024), a general expression for $H_{\rm eff}$} that ensures covariance was established. It takes the form:
${H_{\rm eff}}=-2E^{2}\Big{[}\partial_{s_{1}}M_{\rm eff}+\frac{\partial_{s_{2}}M_{\rm eff}}{2}s_{3}+\frac{\partial_{s_{4}}M_{\rm eff}}{s_{4}}s_{5}+\mathcal{R}\Big{]},$ (1)
where $s_{a}$ for $a=1,2,\cdots,5$ are defined by
$s_{1}=E^{1},\ s_{2}=K_{2},\ s_{3}=\frac{K_{1}}{E^{2}},\ s_{4}=\frac{\partial_{x}s_{1}}{E^{2}},s_{5}=\frac{\partial_{x}s_{4}}{E^{2}},$
$\mathcal{R}$ is an arbitrary function of $s_{1}$ and $M_{\rm eff}$, and the mass function $M_{\rm eff}$ depending on $s_{1},s_{2},s_{4}$ is a solution to the following so-called covariance equation:
$\frac{\mu s_{1}s_{4}}{4}$ $=(\partial_{s_{2}}M_{\rm eff})\partial_{s_{2}}\partial_{s_{4}}M_{\rm eff}-(\partial_{s_{4}}M_{\rm eff})\partial_{s_{2}}^{2}M_{\rm eff},$ (2a)
$0=$ $(\partial_{s_{2}}\mu)\partial_{s_{4}}M_{\rm eff}-(\partial_{s_{2}}M_{\rm eff})\partial_{s_{4}}\mu.$ (2b)
In the present paper, $\mu$ is treated as an auxiliary variable, though it was introduced initially in Zhang et al. (2025, 2024) to deform the classical constraint algebra and encode QG effects.
As in classical GR, the metric is constructed from the phase-space variables together with the freely specifiable Lagrange multipliers $N$ and $N^{x}$, which are respectively known as the lapse function and shift vector. The metric must be defined consistently with $H_{\rm eff}$} so that it remains covariant under different choices of $N$ and $N^{x}$. As shown in Zhang et al. (2025, 2024), the metric compatible with $H_{\rm eff}$} in Eq. (1) is
[…]
in which $\mu$ must be included. Here $\differential\Omega^{2} d$enotes the metric on the unit sphere $\mathbb{S}^{2}$. Note that this metric represents the most general spherically symmetric expression (see, e.g., Fischler et al. (1990)).
In classical GR, the mass function is $M_{\rm cl}=\frac{\sqrt{s_{1}}}{2}\left[1+(s_{2})^{2}-(s_{4})^{2}/4\right]$ which is a solution to Eq. (2) with $\mu=1$. Besides, some other special solutions to the covariance equation were proposed in Zhang et al. (2025, 2024), leading to effective Hamiltonian constraints that depend on a quantum parameter. The classical Hamiltonian constraint is recovered in the limit where this quantum parameter approaches 0. Models based on these special solutions have been extensively studied (see, e.g., Liu et al. (2024); Konoplya and Stashko (2025a); Lin et al. (2025); Belfaqih et al. (2025); Du et al. (2025); Konoplya and Stashko (2025b); Ongole et al. (2025); Motaharfar and Singh (2025); Ai et al. (2025); Del Águila and Morales-Técotl (2025)). Nevertheless, two key open questions remain: how the dynamics behaves when $M_{\rm eff}$ and $\mathcal{R}$ are taken in their most general forms, and what the complete solution space of the covariance equation is. In this letter, we systematically address these two open problems, and show that, for any one-parameter family of static, spherically symmetric spacetime metrics, there exists a corresponding Hamiltonian constraint. Each theory described by the corresponding Hamiltonian constraint takes the given metrics as its unique vacuum solution. In this way, we resolve the inverse problem of Birkhoff’s theorem.
II Dynamics generated by general $H_{\rm eff}$}
As shown in Zhang et al. (2025, 2024), $H_{\rm eff}$} and $H_{x}$ forms a closed algebra under the Poisson bracket, implying that the model is a totally constrained system of the first class. Thus, to solve the dynamics, some gauge fixing condition is necessary. In this paper, we adopt the gauge $E^{1}=x^{2}$ and $E^{2}=\partial_{x}E^{1}/2=x$. In classical GR, this gauge corresponds to choosing the Painlevé–Gullstrand (PG) coordinates, and thus we will refer to it as the PG gauge.
Under the PG gauge, we solve the diffeomorphism constraint to obtain $K_{1}=\partial_{x}K_{2}$. Substituting this result into the effective Hamiltonian constraint in Eq. (1) and vanishing the result, we get $M_{\rm eff}=M(x)$ with $M(x)$ satisfying
$\partial_{x}M(x)+2x\mathcal{R}(x^{2},M(x))=0.$ (4)
From $M_{\rm eff}=M(x)$, we can solve for $s_{2}=K_{2}$ as a function of $x$, since, by definition, $M_{\rm eff}$ depends on the phase space dependent variables $s_{1}$, $s_{2}$, and $s_{4}$, where $s_{1}=x^{2}$ and $s_{4}=2$ are already fixed in the PG gauge. The lapse function $N$ and the shift vector $N^{x}$ are determined by the evolution equations for $E^{I}$, i.e., $\dot{E}^{I}=0=\{E^{I},{H_{\rm eff}}[N]+H_{x}[N^{x}]\},$ yielding (see Appendix V for more details)
[…]
$N^{x}(x)=$ $-\frac{N(x)}{x}(\partial_{s_{2}}M_{\rm eff})(x).$
Here, $(\partial_{M_{\rm eff}}\mathcal{R})(x)$ denotes the partial derivative of $\mathcal{R}(s_{1},M_{\rm eff})$ with respect to $M_{\rm eff}$, evaluated at $s_{1}=x^{2}$ and $M_{\rm eff}=M(x)$. The term $(\partial_{s_{2}}M_{\rm eff})(x)$ is understood analogously. In addition, as shown in Appendix V, we have $\dot{M}_{\rm eff}=0$ under the gauge. As a consequence, the integration constant appearing in the solution to Eq. (4) is a Dirac observable of the model.
According to Eq. (2b), $\mu$ is an arbitrary function of $s_{1}$ and $M_{\rm eff}$. Hence, the dependence of $\mu$ on $x$ can be obtained by substituting $s_{1}\equiv E^{1}=x^{2}$ and $M_{\rm eff}=M(x)$. Applying this result, together with the expressions of $E^{I}$ and Eq. (5), we can obtain the $x$-dependence of the metric in Eq. (3). Obviously, the metric is uniquely determined, with the integration constant, i.e., the Dirac observable, entering through $M(x)$. In this way, we prove the analogue of Birkhoff’s theorem for the models associated with $H_{\rm eff}$}. Since $H_{\rm eff}$} is determined by requiring covariance, this result establishes a connection between Birkhoff’s theorem and general covariance.
III General solution to the covariance equation
To solve Eq. (2a), let us introduce $\tilde{s}_{4}=(s_{4})^{2}$, and $P(s_{1},M_{\rm eff})$ satisfying $\partial_{M_{\rm eff}}P=8/(\mu s_{1})$. Then, Eq. (2a) can be simplified to
$1=(\partial_{s_{2}}P)\partial_{\tilde{s}_{4}}\partial_{s_{2}}M_{\rm eff}-(\partial_{\tilde{s}_{4}}P)\partial_{s_{2}}^{2}M_{\rm eff},$ (6)
which leads to (see Appendix A)
$(\partial_{s_{2}}M_{\rm eff})^{2}=\frac{\mu s_{1}}{4}\left[(s_{4})^{2}+\mathcal{Z}\right],$ (7)
for an arbitrary function $\mathcal{Z}$ of $s_{1}$ and $M_{\rm eff}$. The equivalence between Eq. (2a) and Eq. (7) follows directly from substituting the latter into the former. Inserting Eq. (7) into Eq. (5) and following the procedure described thereafter, we arrive at the explicit expression of the metric under the PG coordinates. It will be convenient to rewrite the result in the Schwarzschild coordinates $(t_{s},x,\theta,\phi)$ as
[…]
where $N$ is given by Eq. (5). The functions $\mu$ and $\mathcal{Z}$, defined as functions of $s_{1}$ and $M_{\rm eff}$, are also to be evaluated at $s_{1}=x^{2}$ and $M_{\rm eff}=M(x)$.
Given any static, spherically symmetric metric, such as the Schwarzschild one, the Hayward one, or those inspired by candidate QG theories, we can identify the corresponding the functions $\mu$, $\mathcal{Z}$, and $\mathcal{R}$ by comparing it with Eq. (8). Then, substituting $\mu$ and $\mathcal{Z}$ into Eq. (7) determines $M_{\rm eff}$ as a function of $s_{1}$, $s_{2}$, and $s_{4}$. This $M_{\rm eff}$, together with the expression of $\mathcal{R}$, can subsequently be inserted into Eq. (1) to obtain $H_{\rm eff}$}. In this way, the generally covariant effective dynamics of the given geometry are reconstructed, as detailed in the next section.
IV reconstructing covariant dynamics from metric
Consider a one-parameter family of static, spherically symmetric spacetimes with the line element
[…]
where $F$ and $H$ are arbitrary functions of $x$, parametrized by the parameter $m$. Comparing Eq. (9) with Eq. (8), we find
$F=-\frac{1}{4}N^{2}\mathcal{Z},\quad H=-\frac{1}{4}\mu\mathcal{Z}.$ (10)
To determine the functions $\mathcal{R}$, $\mu$ and $\mathcal{Z}$ from these two equations, one must first specify one of the functions as input. We choose to input the function $\mathcal{R}(s_{1},M_{\rm eff})$, which allows us to solve Eq. (4) and obtain $M(x)$. Identifying the integration constant arising from the differential equation (4) with $m$, and using the previously established relation $M_{\rm eff}=M(x)$, we can solve for $m$ to get
$m=A(M_{\rm eff},x)$ (11)
which defines a function $A$. Inserting this result into Eq. (10) and applying $s_{1}=x^{2}$, we finally have
$\mu$ $=\frac{N(\sqrt{s_{1}})^{2}H(\sqrt{s_{1}};A(M_{\rm eff},\sqrt{s_{1}}))}{F(\sqrt{s_{1}};A(M_{\rm eff},\sqrt{s_{1}}))},$ (12)
$\mu\mathcal{Z}$ $=-4H(\sqrt{s_{1}};A(M_{\rm eff},\sqrt{s_{1}})).$
Here, $N$ as a function of $\sqrt{s_{1}}=x$ is given by Eq. (5) with $\mathcal{R}$ taken as the previously specified input function. Combining Eqs. (12) and (7), we finally have
$\pm\frac{\sqrt{s_{1}}}{2}s_{2}+\Xi(s_{1},s_{4})$ (13)
[…]
$-4H(\sqrt{s_{1}};A(\xi,\sqrt{s_{1}}))\Bigg{\}}^{-1/2},$
for arbitrary integration constant $\Xi$. From Eq.  (13), we can solve for $M_{\rm eff}$ as a function of $s_{1}$, $s_{2}$ and $s_{4}$ explicitly. Then, substituting it into Eq. (1) yields the expression of $H_{\rm eff}$}, which governs the dynamics of the model.
In what follows, we show explicit forms of $M_{\rm eff}$ for several concrete examples. As the first example, we consider the classical Schwarzschild spacetime with
$H=F=1-\frac{2m}{x}.$ (14)
Choosing $\mathcal{R}=0$ and substituting Eq. (14) into Eq. (13), we finally have
$M_{\rm eff}=\frac{\sqrt{s_{1}}}{2}\left[1+(s_{2})^{2}-\frac{(s_{4})^{2}}{4}\right]+2\frac{\Xi^{2}}{\sqrt{s_{1}}}\pm 2s_{2}\Xi.$ (15)
The mass function $M_{\rm cl}$ in GR can be recovered by choosing $\Xi=0$.
As the second example, we consider the loop quantum black hole spacetime with
$H=F=1-\frac{2m}{x}\left(1-\frac{2\zeta^{2}m}{x^{3}}\right).$ (16)
Note that this metric was previously derived in, e.g., Kelly et al. (2020); Lewandowski et al. (2023), using different approaches, but always in the absence of a covariant dynamical framework. Choosing $\mathcal{R}=0$ and plugging Eq. (16) into Eq. (13), we finally have
$M_{\rm eff}=$ $\frac{\sqrt{s_{1}}^{3}}{2\zeta^{2}}\sin^{2}\left(\frac{\zeta s_{2}}{\sqrt{s_{1}}}\pm\frac{2\zeta\Xi}{s_{1}}\right)$
$\mp\frac{s_{1}\sqrt{(s_{4})^{2}-4}\sin\left(\frac{2\zeta s_{2}}{\sqrt{s_{1}}}\pm\frac{4\zeta\Xi}{s_{1}}\right)}{4\zeta}.$
To ensure the correct classical limit, i.e., that $M_{\rm eff}$ returns to the expression in classical GR as $\zeta$ approaches 0, we should choose $\Xi=\sqrt{s_{1}(s_{4})^{2}-4s_{1}}/4.$
For the Hayward metric Hayward (2006), we have
$F=H\equiv F_{H}=1-\frac{2mx^{2}}{x^{3}+2\zeta^{2}m}.$ (17)
With choosing $\mathcal{R}=0$, we finally have
$-\frac{\left(2\zeta^{2}M_{\rm eff}+\sqrt{s_{1}}^{3}\right)\sqrt{(s_{4})^{2}-4F_{H}(M_{\rm eff},\sqrt{s_{1}})}}{8s_{1}+2\zeta^{2}\left((s_{4})^{2}-4\right)}$
$=$ $\frac{2\sqrt{s_{1}}^{5}\text{arctanh}\left(\frac{\zeta\sqrt{(s_{4})^{2}-4F_{H}(M_{\rm eff},\sqrt{s_{1}})}}{\sqrt{4s_{1}+\zeta^{2}\left((s_{4})^{2}-4\right)}}\right)}{\zeta\sqrt{4s_{1}+\zeta^{2}\left((s_{4})^{2}-4\right)}^{3}}\mp\frac{1}{2}\sqrt{s_{1}}s_{2},$
which defines $M_{\rm eff}$ as a function of $s_{1}$, $s_{2}$ and $s_{4}$ implicitly. In the Hayward case, we have chosen $\Xi=0$ to ensure the classical limit.
V Conclusion and discussion
This work establishes a foundational result: any one-parameter family of static, spherically symmetric metrics can be associated with a family of underlying generally covariant theories in which the metrics arise as the solutions. Remarkably, within each such theory, the given one-parameter family of metrics forms the unique solution space of the theory, i.e., Birkhoff’s theorem in these models remains valid. The freedom in constructing these theories is characterized by the arbitrary functions $\Xi$ and $\mathcal{R}$.
As demonstrated in Zhang et al. (2025, 2024), Eq. (1) follows from three fundamental assumptions: 1) the diffeomorphism constraint remains its classical form; 2) the constraint algebra retains its geometric interpretation as generating hypersurface deformations; 3) the effective Hamiltonian constraint depends only on the same basic scalars $s_{a}$ as in the classical GR. Therefore, given a static metric, the underlying generally covariant theory is uniquely determined, up to the choice of the functions $\Xi$ and $\mathcal{R}$, as long as these physically motivated assumptions are imposed. Moreover, the assumption 3) ensures that each of such theories contains at most second order derivatives of metrics. This complete and robust reconstruction framework establishes the covariant origin of any static, spherically symmetric spacetime, providing a solid foundation for exploring quantum-gravity-inspired and regular black hole geometries.
Acknowledgements.
This work was supported in part by the National Key Research and Development Program of China Grant No. 2021YFC2203001 and in part by the NSFC (Grants No. 12275022, No. 12475046, and No. 12505055) and “the Fundamental Research Funds for the Central Universities”.
Appendices
Appendix A Solving the dynamics
Under the PG gauge, the evolution equation $\dot{E}^{I}=0$ leads to
$0=$ $-\frac{2N}{x}\frac{\partial{H_{\rm eff}}}{\partial s_{3}}+2xN^{x},$ (A1)
$0=$ $-N\frac{\partial{H_{\rm eff}}}{\partial s_{2}}+\partial_{x}(xN^{x}).$
Due to Eq. (1), we have
$\frac{\partial{H_{\rm eff}}}{\partial s_{2}}=$ $-\partial_{x}(\partial_{s_{2}}M_{\rm eff})-2x\partial_{M_{\rm eff}}\mathcal{R}\partial_{s_{2}}M_{\rm eff},$
$\frac{\partial{H_{\rm eff}}}{\partial s_{3}}=$ $-x\partial_{s_{2}}M_{\rm eff},$
where the diffeomorphism constraint $H_{x}=0$ is applied for the first equation. Thus, Eq. (A1) can be simplified to be
$0=$ $N\partial_{s_{2}}M_{\rm eff}+xN^{x},$
$0=$ $N\partial_{x}(\partial_{s_{2}}M_{\rm eff})+2Nx\partial_{M_{\rm eff}}\mathcal{R}\partial_{s_{2}}M_{\rm eff}+\partial_{x}(xN^{x}),$
which finally leads to Eq. (5). For $\dot{M}_{\rm eff}$, a straightforward calculation shows
$\dot{M}_{\rm eff}=-\frac{N\partial_{s_{2}}M_{\rm eff}}{E^{2}}\left(2E^{2}\mathcal{R}+\partial_{x}M_{\rm eff}\right)=0,$
where $E^{2}=\partial_{x}E^{1}/2$ as a result of the PG gauge and the first equation in Eq. (A1) are applied for the first equality, and the PG gauge and Eq. (4) are applied to get the last equality.
Appendix B Solving the covariance equation
Let us start from Eq. (6). By defining $Q=\partial_{s_{2}}M_{\rm eff}$, we can rewrite it as
$\det\begin{pmatrix}\partial_{s_{2}}P&\partial_{\tilde{s}_{4}}P\\ \partial_{s_{2}}Q&\partial_{\tilde{s}_{4}}Q\end{pmatrix}=1.$
This result, combined with the properties of the Jacobian matrix, leads to
$\begin{pmatrix}\partial_{P}s_{2}&\partial_{Q}{s_{2}}\\ \partial_{P}{\tilde{s}_{4}}&\partial_{Q}{\tilde{s}_{4}}\end{pmatrix}=\begin{pmatrix}\partial_{\tilde{s}_{4}}Q&-\partial_{\tilde{s}_{4}}P\\ -\partial_{s_{2}}Q&\partial_{s_{2}}P\end{pmatrix}$.
Then, we have
$\frac{\partial\tilde{s}_{4}}{\partial Q}=\partial_{s_{2}}P=Q\partial_{M_{\rm eff}}P,$
leading to Eq. (7) ultimately.